# Towards a universal QAOA protocol: Evidence of a scaling advantage in solving some combinatorial optimization problems

### v2

@@ Abstract
The quantum approximate optimization algorithm (QAOA) is a promising algorithm for solving combinatorial optimization problems (COPs). In this algorithm, there are alternating layers consisting of a mixer and a problem Hamiltonian. Each layer $i=0,\ldots,p-1$ is parameterized by $\beta_{i}$ and $\gamma_{i}$. How to find these parameters has been an open question with the majority of the research focused on finding them using classical algorithms. In this work, we present evidence that fixed linear ramp schedules constitute a universal set of QAOA parameters, i.e., a set of $\gamma$ and $\beta$ parameters that rapidly approximate the optimal solution, $x^{*}$, independently of the COP selected, and that the success probability of finding it, $probability(x^{*})$, increases with the number of QAOA layers $p$. We simulate linear ramp QAOA protocols (LR-QAOA) involving up to $N_{q}=42$ qubits and $p=400$ layers on random instances of 9 different COPs. The results suggest that $probability(x^{*})\approx 1/2^{(\eta N_{q}/p)}$ for a constant $\eta$. We extend the analysis in 4 COPs with $p=N_{q}$ and show that $probability(x^{*})$ seems to be constant for general cases. For example, when implementing LR-QAOA with $p=42$, the $probability(x^{*})$ for 42-qubit Weighted MaxCut problems (W-MaxCut) increases from $2/2^{42}\approx 10^{-13}$ to an average of 0.13. We compare LR-QAOA, simulated annealing (SA), and branch-and-bound (B&B) finding a scaling improvement in LR-QAOA. We test LR-QAOA on real hardware using IonQ Aria, Quantinuum H2-1, IBM Brisbane, IBM Kyoto, and IBM Osaka, encoding random weighted MaxCut (W-MaxCut) problems from 5 to 109 qubits and $p=3$ to $100$. We find that even for the largest case, $N_{q}=109$ qubits and $p=100$, information about the LR-QAOA optimization protocol is still present. The circuit involved requires 21200 CNOT gates and a time of $\approx 132\ \mu s$. The resilience of LR-QAOA to noise is attributed to its characteristic of bringing the system towards a minimal energy state despite noise driving it towards a maximally mixed state. These results show that LR-QAOA effectively finds high-quality solutions for a large variety of COPs and suggest a scaling advantage of quantum computation for combinatorial optimization.
@@ I Introduction
Finding high-quality solutions for COPs is perceived as one of the main applications of quantum computation in the near future. In the gate-based regime, QAOA [1] has become one of the most studied quantum algorithms for solving COPs. There are different factors for the extensive study of QAOA. Firstly, parametric unitary gates can effectively represent the Hamiltonian of the COPs, where the ground state encodes the optimal solution of the problem [2, 3]. Moreover, QAOA has a performance guarantee in the limit of infinite layers resembling the quantum adiabatic algorithm [4, 1]. Additionally, QAOA needs fewer resources (e.g. number of gates and qubits) compared to other quantum algorithms and can be tested on current state-of-the-art quantum hardware [5, 6, 7]. Furthermore, classical methods find it hard to solve large instances of COPs with practical applications [8], and therefore finding alternative ways to solve them is needed. Ultimately, the goal of quantum optimization algorithms, as exemplified by QAOA, is to demonstrate advantages in solving optimization problems, be it in terms of energy efficiency, time-to-solution (TTS), or solution quality compared to classical methods.
In the simplest version of QAOA, the cost Hamiltonian of a combinatorial optimization problem is encoded in a parametric unitary gate along with a “mixer”, a second parametric unitary gate that does not commute with the first unitary gate. In this context, parameters $\gamma=[\gamma_{0},...,\gamma_{p-1}]$ and $\beta=[\beta_{0},...,\beta_{p-1}]$ for the cost and mixer Hamiltonians, respectively, are adjusted to minimize the expectation value of the cost Hamiltonian for $i=0,...,p-1$ layers of QAOA. Since its conception, a classical algorithm was suggested to find the QAOA $\gamma$ and $\beta$ parameters [1]. This makes QAOA fall in the category of variational quantum algorithms (VQA) [9]. However, these algorithms have exhibited a limited/poor performance advantage as the classical optimization part finds it hard to escape local minima when searching for $\gamma$ and $\beta$ parameters [10, 11].
Modest progress has been made by thinking of QAOA as a VQA, with major studies conducted in regimes of a few qubits and shallow circuits [5]. Deep QAOA circuits, when viewed through the lens of VQA, lead to a pessimistic conclusion regarding their universal applicability [12]. Moreover, implementations on real hardware face an even greater challenge, the noise inherent in current quantum devices makes the search for the minima of the objective function unfeasible after only a few QAOA layers [13, 12].
@@
Various protocols have been proposed to fix these parameters. In [11], Krementski et. al. found a set of QAOA parameters with consistent performance to find optimal solutions using a fixed LR-QAOA protocol. They tested this methodology using the Hamiltonian of different molecules, an Ising Hamiltonian, and the 3-SAT problem for intermediate to large $p$. Another attempt for fixing the QAOA protocol is proposed in [17], in which the authors presented QAOA as a second-order time discretization of QA referred to as approximate quantum annealing (AQA). LR-QAOA can be considered as an AQA protocol with a linear annealing schedule.
@@
In [18], we proposed fixed schedules transferring optimal $\gamma$ and $\beta$ parameters between different COPs. We found that $\gamma$ and $\beta$ parameters that work well for a COP eventually give good results in other COPs independently of the COP structure. This information led us to the results of the present work.
@@
Recently, Kremenetski et. al. have explained the behavior of LR-QAOA and in general of the gradually changing schedules using the discrete adiabatic theorem involving a wrap-around phenomenon [19].
@@
In this paper, we extend the study of LR-QAOA schedules to different COPs presenting numerical and experimental evidence that LR-QAOA constitutes a universal QAOA protocol, i.e., the set of $\gamma$ and $\beta$ parameters from a linear ramp schedule that work effectively independently of the problem or problem size in combinatorial optimization. We test this protocol using maximal independent set (MIS), bin packing (BPP), traveling salesman (TSP), maximum cut (MaxCut), weighted maximum cut (W-MaxCut), 3-regular graph maximum cut (3-MaxCut), Knapsack (KP), portfolio optimization (PO), maximum 2 Boolean satisfiability problem (Max-2-SAT), and maximum 3-SAT (Max-3-SAT). We use random instances of these COPs with problem sizes ranging from 4 to 42 qubits and $p$ from 3 to 400. For large problems, we simulate them using JUQCS–G software [17] on JUWELS Booster, a cluster of 3744 NVIDIA A100 Tensor Core GPUs, integrated into the modular supercomputer JUWELS [20, 21].
@@
In these cases, the average probability of success over the 5 random instances seems to follow a scaling that can be described by $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for a constant $\eta$. We extend the analysis on 4 different problems keeping the relation $p=N_{q}$, expecting $probability(x^{*})$ to stay constant. We use fully connected random W-MaxCut, MaxCut, Max-2-SAT, and Max-3-SAT. The W-MaxCut, MaxCut, and Max-3-SAT are both NP-Hard [22] and APX-Hard [23] problems, and we show that the time-to-solution (TTS) [13] using our selected instances of W-MaxCut indeed grows exponentially when using classical optimizers.
We find a scaling improvement in terms of the TTS when using LR-QAOA with the probability of success on average remaining constant $probability(x^{*})=1/2^{\eta}$ using $p=N_{q}$. With this further evidence, we conjecture that the expected probability of success will follow the relation $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for typical cases of COPs. If this conjecture holds, it implies that there is a quantum algorithm that in polynomial time can solve instances of NP-Hard problems that are already hard for classical algorithms. This linear time complexity has also been observed recently for 3-MaxCut problems in [24] using what they called Floquet adiabatic algorithm which follows the same ansatz that LR-QAOA but from the point of view of a Trotterization approximation of the adiabatic algorithm similar to AQA. However, we note that for some cases, the relation $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ seems not to hold. We attributed it to the high concentration of sub-optimal solution close to the optimal. Even if the conjecture does not hold, there is a noticeable scaling advantage in LR-QAOA compared to SA and B&B for solving instances of W-MaxCut. This evidence complements the recent findings in [25] where a scaling advantage is observed of a fixed QAOA protocol for solving a classical intractable COP known as low autocorrelation binary sequences (LABS).
@@
We extend the analysis to real quantum hardware. Using IonQ Aria (ionq_aria), Quantinuum H2-1 (quantinuum_H2) [26], IBM Brisbane (ibm_brisbane), IBM Kyoto (ibm_kyoto), and IBM Osaka (ibm_osaka), we run W-MaxCut problems ranging from 5 to 109 qubits and $p$ from 3 to 100. We find that there is an effective number of layers, $p_{\mathrm{eff}}$, for which the best performance is obtained using each device. In the case of IBM devices $p_{\mathrm{eff}}=10$, on ionq_aria $p_{\mathrm{eff}}=10$, and on quantinuum_H2 $p_{\mathrm{eff}}=50$. Remarkably, for the largest problem size, 109 qubits and $p=100$, we observe that LR-QAOA still possesses an advantage over random sampling in ibm_kyoto and ibm_osaka. For a comparative analysis between the different vendors, we test a 25-qubit W-MaxCut problem on them, quantinuum_H2 gives the best performance with a probability($x^{*}$) $=0.08$ at $p=50$.
We demonstrate that LR-QAOA is resilient to noise. This robustness is attributed to the interplay of two opposing forces. While noise pushes the system towards a maximally mixed state, LR-QAOA drives the system towards the minimum energy of the cost Hamiltonian correcting errors on the way. To explore this phenomenon, we conduct simulations using a Qiskit Aer noise model [27] on a 10-qubit W-MaxCut problem. We vary the strengths of depolarizing noise affecting the 2-qubit gates within the algorithm. Our findings show significant alignment with the behavior exhibited by ionq_aria when solving the same problem.
The rest of the paper is organized as follows. Section II provides a description of the LR-QAOA protocol, its universal properties, a description of the COPs used, the classical solvers used for comparison, the Hamiltonian normalization, the performance diagram, the mitigation technique, the experimental details, and the noise simulation. In Sec. III, we present our results and a discussion for simulation and experimentation using the given protocol. Finally, Sec. IV contains our conclusions.
@@ II.1 LR-QAOA
where $\gamma_{i}$ is a parameter that in our case comes from the linear ramp schedule. Following this, in every second part of a layer, a unitary operator is applied given by
@@
for $i=0,...,p-1$. For our simulations and experimental results, we use $\Delta_{\beta}=0.3$ and $\Delta_{\gamma}=0.6$. In the extended simulations for the conjecture, we use $\Delta_{\beta}=\Delta_{\gamma}=\Delta_{\gamma,\beta}=0.5$ and $\Delta_{\gamma,\beta}=0.4$.
Figure 1: (a) Quantum circuit of the QAOA algorithm, (b) LR-QAOA schedule, (c) eigenvalue distribution for the different COPs. All the distributions are for 10-qubit problems except BPP and TSP with 12-qubit and 9-qubit respectively. (d) An optimal solution for one of the 42-qubit W-MaxCut problems using $p=50$ LR-QAOA. Dashed lines represent cuts, black (white) vertices qubits in 1 (0) state. At the end of the algorithm, the probability of finding the maximum cut is 32%.
II.2 Universal properties of LR-QAOA
@@
where $t$ is some step in the QAOA algorithm, $k$ are the states in the computational basis, and $\alpha_{k}^{t}$ the amplitude of $|k\rangle$ at time $t$.
The unitary transformation induced by $U_{C}(\gamma_{t})$ produces a rotation in the complex plane for every state given by
$\alpha_{k}^{t+1}=e^{i\theta_{k}^{t}}\alpha_{k}^{t},$ (6)
@@
where $E_{k}=\langle k|H_{c}|k\rangle$. This evolution is shown in Fig. 2 (a). Equation 7 explains why the amplitude amplification is proportional to the energy of a given solution. Negative energies are rotated counterclockwise with the rotation proportional to their energies. This can be seen in Fig. 2-(e).
The change by $U_{B}(\beta_{t})$ is more complex and depends on the Hamming distance of the given state to the other states. This operator is responsible for the change in energy and produces an interference pattern that exploits the $U_{C}(\gamma_{t})$ effect. It is described by
$\alpha^{t+1}_{k}=\sum_{l=0}^{2^{N}_{q}-1}\left(\cos(\beta_{t})\right)^{N_{q}-k\cdot l}\left(j\sin(\beta_{i})\right)^{k\cdot l}\alpha^{t}_{l},$ (8)
@@
$k\cdot l=\sum_{m=0}^{N_{q}-1}(k[m]+l[m])\mod 2,$ (9)
where $k$ and $l$ are states in the computational basis. Equation 9 gives the Hamming distance between the state $k$ and $l$. For example, in a 3-qubit system, the evolution of $\alpha^{t}_{000}$ is given by
@@
In Fig. 2-(c) is shown the evolution of the optimal solution, $x^{*}$, of a 6-qubit W-MaxCut problem for the $U_{C}(\gamma_{t})$ and $U_{B}(\beta_{t})$ steps for $t\in\{0,...,Nq-1\}$. In Fig.2-(d) shows the same evolution but for the state with the lowest energy. In this case, the evolution of the eigenvalue due to $U_{C}(\gamma_{t})$ goes in the opposite direction to the evolution of $U_{B}(\beta_{t})$ producing the desired effect of interference. Fig.2-(e) shows the angle of rotation of the white triangles, i.e., the rotation due to $\theta(\gamma_{t})$. The gain in the amplitude of the $\alpha_{k}^{t}$ at each time step after the unitary evolution $U_{B}(\beta_{t})$ is shown in Fig. 2-(f).
Figure 2: LR-QAOA from the point of view of individual time steps. (a) The action of the $U_{C}(\gamma_{t})$ gate on the state $k$ at time step $t$, (b) evolution of the $|000\rangle$ amplitude after the application of $U_{B}(\beta_{t})$ for a 3-qubit system, (c) evolution of the optimal solution, $x^{*}$, in a 6-qubit W-MaxCut problem. The gray (white) triangles are a time-step evolution due to $U_{B}(\gamma_{t})$ ($U_{C}(\gamma_{t})$). Line colors represent the time steps being blue (red) $t=0$ ($t=5$) step. (d) Evolution of the worst solution for the 6-qubit W-MaxCut problem. (e)-left LR-QAOA $\gamma$ rotations at each layer for each state. Positive angles refer to counterclockwise rotations. Colors represent the energy of the state with darker blue (red) closer to the optimal (worst) solution of the problem. (e)-right last layer rotation in LR-QAOA vs. the energy, following Eq 7. (f) Amplitude gain evolution of the states after each $U_{B}(\beta_{t})$ for the 6-qubit W-MaxCut problem.

### v3

# Towards a Linear-Ramp QAOA protocol: Evidence of a scaling advantage in solving some combinatorial optimization problems

@@ Abstract
The Quantum Approximate Optimization Algorithm (QAOA) is a promising algorithm for solving combinatorial optimization problems (COPs), with performance governed by variational parameters $\{\gamma_{i},\beta_{i}\}_{i=0}^{p-1}$. While most prior work has focused on classically optimizing these parameters, we demonstrate that fixed linear ramp schedules, linear ramp QAOA (LR-QAOA), can efficiently approximate optimal solutions across diverse COPs. Simulations with up to $N_{q}=42$ qubits and $p=400$ layers suggest that the success probability scales as $P(x^{*})\approx 2^{-\eta(p)N_{q}+C}$, where $\eta(p)$ decreases with increasing $p$. For example, in Weighted Maxcut instances, $\eta(10)=0.22$ improves to $\eta(100)=0.05$. Comparisons with classical algorithms, including simulated annealing, Tabu Search, and branch-and-bound, show a scaling advantage for LR-QAOA. We show results of LR-QAOA on multiple QPUs (IonQ, Quantinuum, IBM) with up to $N_{q}=109$ qubits, $p=100$, and circuits requiring 21,200 CNOT gates. Finally, we present a noise model based on two-qubit gate counts that accurately reproduces the experimental behavior of LR-QAOA.
@@ I Introduction
Finding high-quality solutions for COPs is perceived as one of the main applications of quantum computation in the near future. In the gate-based regime, QAOA [1] has become one of the most studied quantum algorithms for solving COPs. There are different factors for the extensive study of QAOA. Firstly, parametric unitary gates can effectively represent the Hamiltonian of the COPs, where the ground state encodes the optimal solution of the problem [2, 3]. Moreover, QAOA has a performance guarantee in the limit of infinite layers resembling the quantum adiabatic algorithm [4, 1]. Additionally, QAOA needs fewer resources (e.g., number of gates and qubits) compared to other quantum algorithms and can be tested on current state-of-the-art quantum hardware [5, 6, 7]. Furthermore, classical methods find it hard to solve large instances of COPs with practical applications [8], and therefore, finding alternative ways to solve them is needed. Ultimately, the goal of quantum optimization algorithms, as exemplified by QAOA, is to demonstrate advantages in solving optimization problems, be it in terms of energy efficiency, time-to-solution (TTS), or solution quality compared to classical methods.
In the simplest version of QAOA, the cost Hamiltonian of a combinatorial optimization problem is encoded in a parametric unitary gate along with a “mixer”, a second parametric unitary gate that does not commute with the first unitary gate. In this context, parameters $\gamma=[\gamma_{0},...,\gamma_{p-1}]$ and $\beta=[\beta_{0},...,\beta_{p-1}]$ for the cost and mixer Hamiltonians, respectively, are adjusted to minimize the expectation value of the cost Hamiltonian for $i=0,...,p-1$ layers of QAOA. Since its conception, a classical algorithm was suggested to find the QAOA $\gamma$ and $\beta$ parameters [1]. This makes QAOA fall in the category of variational quantum algorithms (VQA) [9]. However, these algorithms have exhibited a limited/poor performance advantage as the classical optimization part finds it hard to escape local minima when searching for $\gamma$ and $\beta$ parameters [10, 11]. The barren plateau is another challenge in QAOA. It refers to regions in the cost function landscape where the gradient is nearly zero, making it hard to find QAOA parameters via gradient-based optimization [12].
Modest progress has been made by considering QAOA as a VQA, with major studies conducted in regimes of a few qubits and shallow circuits [5]. Deep QAOA circuits, when viewed through the lens of VQA, lead to a pessimistic conclusion regarding their universal applicability [13]. Moreover, implementations on real hardware face an even greater challenge; the noise inherent in current quantum devices makes the search for the minima of the objective function unfeasible after only a few QAOA layers [14, 13].
@@
Various protocols have been proposed to fix these parameters. In [11], Krementski et. al. found a set of QAOA parameters with consistent performance to find optimal solutions using a fixed LR-QAOA protocol. They tested this methodology using the Hamiltonian of different molecules, an Ising Hamiltonian, and the 3-SAT problem for intermediate to large $p$. Another attempt to fix the QAOA protocol is proposed in [18], in which the authors presented QAOA as a second-order time discretization of QA referred to as approximate quantum annealing (AQA). In [19] is proposed the Trotterized adiabatic evolution (TAE), an idea similar to AQA but using a fixed sinusoidal schedule. LR-QAOA can be considered as an AQA protocol with a linear annealing schedule.
@@
In [20], we proposed fixed schedules transferring optimal $\gamma$ and $\beta$ parameters between different COPs. We found that sometimes $\gamma$ and $\beta$ parameters that work well for a COP in the form of Eq. 1 give good results on other COPs with different structures. Specifically, we found that parameters optimized for the bin packing problem (BPP) can be translated to Maxcut, Maximal independent set (MIS), portfolio optimization (PO), and travelling salesman problem (TSP), giving a quadratic speedup over random guessing on all of them. This suggests that there are effective QAOA protocols that work for different problems. This information led us to the results of the present work.
@@
Recently, Kremenetski et. al. have explained the behavior of LR-QAOA and, in general, of the gradually changing schedules using the discrete adiabatic theorem involving a wrap-around phenomenon [21].
@@
In this paper, we extend the study of LR-QAOA schedules to different COPs, presenting numerical and experimental evidence that LR-QAOA constitutes an effective QAOA protocol, i.e., the set of $\gamma$ and $\beta$ parameters from a linear ramp schedule works effectively for many problems and problem sizes in combinatorial optimization. We test this protocol using MIS, BPP, TSP, Maxcut, weighted maximum cut (WMaxcut), 3-regular graph maximum cut (3-Maxcut), Knapsack (KP), PO, maximum 2 Boolean satisfiability problem (Max-2-SAT), and maximum 3-SAT (Max-3-SAT). We use random instances of these COPs with problem sizes ranging from 4 to 42 qubits and $p$ from 3 to 400. For large problems, we simulate them using JUQCS–G software [18] on JUWELS Booster, a cluster of 3744 NVIDIA A100 Tensor Core GPUs, integrated into the modular supercomputer JUWELS [22, 23].
@@
In these cases, the average probability of success over the 100 random instances seems to follow a scaling that can be described by $probability(x^{*})=2^{-\eta(p)N_{q}+C}$ for a $\eta(p)$ decreasing with $p$ and a constant $C$. We extend the analysis to fully connected random WMaxcut. The WMaxcut is both NP-Hard [24] and APX-Hard [25] problems.
We find a scaling improvement in terms of the time-to-solution (TTS) [14] when using LR-QAOA compared to SA, TABU search, and B&B for solving random instances of WMaxcut. This evidence complements the recent findings in [26] where a scaling advantage is observed in a fixed QAOA protocol for solving a classical intractable COP known as low autocorrelation binary sequences (LABS) and on k-SAT problems [27].
@@
We extend the analysis to real quantum hardware. Using IonQ Aria (ionq_aria), Quantinuum H2-1 (quantinuum_H2) [28], IBM Brisbane (ibm_brisbane), IBM Kyoto (ibm_kyoto), IBM Osaka (ibm_osaka), and IBM fez (ibm_fez), we run WMaxcut problems ranging from 5 to 109 qubits and $p$ from 3 to 100. We find that there is an effective number of layers, $p_{\mathrm{eff}}$, for which the best performance is obtained using each device.
[…]
We present a noise model of LR-QAOA that fits depolarizing noise simulation and experiments on ibm_fez and an emulator of Quantinuum H1-1. The noise model depends only on the number of 2-qubit gates and a noise parameter associated with the QPU. We observe that there is an interplay between the noise pushing the system towards a maximally mixed state and LR-QAOA driving the system towards the minimum energy of the cost Hamiltonian.
@@ II Methods
In this section, we describe the LR-QAOA, some properties of LR-QAOA, the combinatorial optimization problems used, the classical solvers used to compare scaling properties, and experimental details on real quantum hardware.
@@ II.1 LR-QAOA
Figure 1: (a) Quantum circuit of the QAOA algorithm, (b) LR-QAOA schedule, (c) Density vs the normalized eigenvalue distribution for the different COPs with $e$ representing the eigenvalue. All the distributions are for 10-qubit problems except BPP and TSP, with 12-qubit and 9-qubit problems, respectively. (d) An optimal solution for one of the 42-qubit WMaxcut problems using $p=50$ LR-QAOA. Dashed lines represent cuts, black (white) vertices qubits in 1 (0) state. At the end of the algorithm, the probability of finding the maximum cut is 32%.
@@
where $\gamma_{i}$ is a parameter that in our case comes from the linear ramp schedule. Following this, in every second part of a layer, a unitary operator is applied, given by
@@
for $i=0,...,p-1$. For our simulations, we scan over a set of $\Delta_{\gamma}$ and $\Delta_{\beta}$ from one problem at each problem size and use the best value over the remaining cases. For the experimental results, we use $\Delta_{\beta}=0.3$ and $\Delta_{\gamma}=0.6$.
II.2 Properties of LR-QAOA
Figure 2: LR-QAOA from the point of view of individual time steps. (a) The action of the $U_{C}(\gamma_{t})$ gate on the state $k$ at time step $t$, (b) evolution of the |000|000$\rangle$| 000 ⟩ amplitude after the application of $U_{B}(\beta_{t})$ for a 3-qubit system, (c) evolution of the optimal solution, $x^{*}$, in a 6-qubit WMaxcut problem. The gray (white) triangles are a time-step evolution due to $U_{B}(\gamma_{t})$ ($U_{C}(\gamma_{t})$). Line colors represent the time steps being blue (red) $t=0$ ($t=5$) step. (d) Evolution of the worst solution for the 6-qubit WMaxcut problem. (e)-left LR-QAOA $\gamma$ rotations at each layer for each state. Positive angles refer to counterclockwise rotations. Colors represent the energy of the state, with darker blue (red) closer to the optimal (worst) solution of the problem. (e)-Right last layer rotation in LR-QAOA vs. the energy, following Eq 7. (f) Amplitude gain evolution of the states after each $U_{B}(\beta_{t})$ for the 6-qubit WMaxcut problem.
@@ II.2 Properties of LR-QAOA
where $t$ is some step in the QAOA algorithm, $k$ is the state in the computational basis, and $\alpha_{k}^{t}$ the amplitude of $|k\rangle$ at time $t$.
The unitary transformation induced by $U_{C}(\gamma_{t})$, $|\psi_{t+1}\rangle=U_{C}(\gamma_{t})|\psi_{t}\rangle$, produces a rotation in the complex plane for every state given by
$\alpha_{k}^{t+1}=e^{j\theta_{k}^{t}}\alpha_{k}^{t},$ (6)
@@
where $E_{k}=\langle k|H_{c}|k\rangle$. This evolution is shown in Fig. 2 (a). Eq. 7 explains why the amplitude amplification is proportional to the energy of a given solution. Negative energies are rotated counterclockwise with the rotation proportional to their energies. This can be seen in Fig. 2-(e).
The change by $U_{B}(\beta_{t})$, $|\psi_{t+1}\rangle=U_{B}(\beta_{t})|\psi_{t}\rangle$, is more complex and depends on the Hamming distance of the given state to the other states. This operator is responsible for the change in energy and produces an interference pattern that exploits the $U_{C}(\gamma_{t})$ effect. It is described by
$\alpha^{t+1}_{k}=\sum_{l=0}^{2^{N_{q}-1}}\left(\cos(\beta_{t})\right)^{N_{q}-k\cdot l}\left(j\sin(\beta_{t})\right)^{k\cdot l}\alpha^{t}_{l},$ (8)
@@
$k\cdot l=\sum_{m=0}^{N_{q}-1}(k_{m}\oplus l_{m}),$ (9)
where $k$ and $l$ are states in the computational basis. Equation 9 gives the Hamming distance between the states $k$ and $l$. See Supplementary Note 5 for a detailed derivation of Eq. 8. For example, in a 3-qubit system, the evolution of $\alpha^{t}_{000}$ is given by
@@
In Fig. 2-(c) is shown the evolution of the optimal solution, $x^{*}$, of a 6-qubit WMaxcut problem for the $U_{C}(\gamma_{t})$ and $U_{B}(\beta_{t})$ steps for $t\in\{0,...,Nq-1\}$. In Fig.2-(d) shows the same evolution but for the state with the lowest energy. In this case, the evolution of the eigenvalue due to $U_{C}(\gamma_{t})$ goes in the opposite direction to the evolution of $U_{B}(\beta_{t})$ producing the desired effect of interference. Fig.2-(e) shows the angle of rotation of the white triangles, i.e., the rotation due to $\theta(\gamma_{t})$. The gain in the amplitude of the $\alpha_{k}^{t}$ at each time step after the unitary evolution $U_{B}(\beta_{t})$ is shown in Fig. 2-(f).